# TOMATOES: Topology and Material Optimization for Latent Heat Thermal Energy Storage Devices

### v1

@@ 1 Introduction
The imperative for decarbonization [1], grid stabilization [2], and waste-heat recovery [3] has increased the demand for advanced energy storage technologies. Among competing technologies, latent-heat thermal energy storage (LHTES)—which stores energy in the latent heat of phase-change materials (PCMs)—has emerged as a promising candidate due to its high storage density [4]. The realization of next-generation LHTES systems is seen as critical, supported by advances in both materials and system design.
@@
On the materials front, progress in discovery and design of materials has yielded an expanded palette of energy storage and transport materials with improved storage capacity, thermal conductivity, and stability [5, 4]. In parallel, new manufacturing methods—particularly additive and hybrid manufacturing—have enabled the fabrication of complex heat-conduit geometries that can substantially enhance thermal transport within LHTES systems [6]. Together, these developments create both opportunity and complexity for the practicing engineer in designing optimal LHTES systems.
@@
In particular, effectively leveraging this expanded material and geometric design space to extract optimal performing configurations presents a central challenge [7]. Designers must select from a vast and diverse materials space while accounting for the system’s nonlinear, transient, and strongly coupled thermophysical behavior. Furthermore, the geometric freedom afforded by modern manufacturing surpasses the capabilities of traditional, intuition-driven design workflows. Consequently, these manual design approaches are limited in scope and frequently overlook high-performance configurations.
@@
To overcome these limitations, computational methods offer a systematic path forward. Specifically, topology optimization (TO) [8, 9] and machine learning (ML) [10] have emerged as powerful tools for automating design exploration. TO employs gradient-based optimization to navigate complex design landscapes, systematically discovering novel geometries and material layouts to improve performance while respecting governing physics and design constraints. Complementing this physics-driven approach, ML methods excel at discovering, organizing, and navigating expansive material databases to inform design selection [7].
@@
The integration of TO and ML is thus well-positioned to advance LHTES device performance. A critical step toward this goal is addressing the strong coupling between geometry and material, which necessitates their simultaneous optimization. Hitherto, applications of TO have largely ignored this coupling, focusing primarily on geometric optimization while fixing the material choice a priori [11, 12, 13]. For instance, consider a vertical tube carrying a cold heat-transfer fluid (HTF), which is surrounded by both a low-conductivity PCM and a high-conductivity material (HCM) as shown in fig.˜1(a). In traditional approaches, the PCM and HCM are first selected a priori from a material database (fig.˜1 (b)), and their spatial distribution (fig.˜1(c)) is subsequently optimized using the prescribed material properties. Consequently, the more challenging problem of co-designing geometry and material has received far less attention.
@@
A primary barrier to this simultaneous optimization is that material selection is an inherently discrete problem, whereas gradient-based topology optimization requires a differentiable design space. This fundamental incompatibility between discrete material choice and continuous geometric optimization makes true co-design—the joint optimization of material and geometry—exceptionally challenging.
@@
This paper directly addresses this methodological gap. Building upon recent work [7], we introduce an ML-enabled framework that bridges the discrete and continuous domains. Our approach employs a variational autoencoder (VAE) [14] to embed discrete material candidates into a continuous and differentiable latent space. This learned material representation is then integrated directly into a density-based topology optimization formulation, enabling the simultaneous co-design of material and geometry within a single, unified framework. We demonstrate that this co-design methodology yields LHTES devices with performance superior to both conventional baseline designs and sequential optimization workflows.
@@ 2 Related Work
This section surveys the state of the art in the design of LHTES devices. The survey is structured around the central contribution of this paper: the concurrent material selection and topology optimization through an ML augmented framework for the design of LHTES devices. We organize the review into three subsections. The first subsection reviews TES devices and the materials commonly employed therein. The second subsection surveys the use of TO in thermal management; in particular LHTES devices. The third subsection examines the co-design of material and geometry, outlining the limitations of conventional approaches and recent developments that incorporate machine learning.
@@ 2.1 Thermal Energy Storage
The rapid growth of renewable energy and the increasing variability of supply and demand highlight the need for next generation energy storage systems [15]. A wide range of storage technologies has been developed to address these challenges, including electrochemical systems [16], mechanical storage [17], and TES [15], each with distinct operating principles, efficiency levels, costs, and environmental impacts [18, 19]. Among these, TES plays an important role in applications involving heat management and renewable energy integration [20]. TES systems are typically classified into three categories: sensible heat, latent heat, and thermochemical storage [15]. Sensible storage, while widely deployed in hot water tanks and packed-bed systems, generally exhibit lower storage densities compared with latent and thermochemical TES types [15, 21]. Furthermore, thermochemical storage can achieve high energy density, but it relies on reversible chemical reactions that require high material stability and introduce significant complexity in terms of kinetics and long-term durability [22, 18]. Whereas, latent heat storage combines high storage density with nearly isothermal charging and discharging, which makes it well suited for renewable integration and thermal management applications [15, 12, 23].
LHTES systems rely on PCMs as the medium for storing and releasing energy [11]. The performance of an LHTES device therefore depends critically on the type of PCM employed and how it is integrated with HCMs for heat transfer. Broadly, PCMs are categorized into three groups: organics, inorganics, and eutectics [15]. To begin with, organic PCMs , such as paraffins and fatty acids, are widely used because they are chemically stable, non-corrosive, and readily available [15]. However, their low thermal conductivity often limits heat transfer rates, which necessitates the inclusion of HCMs such as copper or aluminum to improve performance [24, 25, 12]. In contrast, inorganic PCMs, particularly salt hydrates, offer higher volumetric storage capacity and thermal conductivity. Yet, they are frequently affected by drawbacks such as supercooling, phase segregation, and material compatibility issues that hinder long-term reliability [15]. Alternatively, eutectic PCMs combine two or more components in fixed ratios, producing well-defined melting points tailored to specific operating conditions [15]. This makes them suitable for applications where precise temperature control is required. This diversity of PCM options and associated trade-offs motivates the need for advanced design approaches; the next subsection reviews the use of TO as a methodology for designing efficient LHTES systems.
@@ 2.2 Topology Optimization of Thermal Energy Storage Devices
TO has transformed the design of structural [8], fluid [26, 27, 28], photonics [29, 30], and thermal systems [31], enabling automated generation of high-performance geometries that outperform traditional, manually designed configurations. Its relevance has increased further with the advent of additive manufacturing, which enables the fabrication of the complex, non-intuitive geometries produced through optimization studies [32]. In thermal management applications, TO has been employed to enhance the performance of heat sinks, heat exchangers, and cooling plates, producing designs that consistently outperform conventional fin-and-tube solutions in terms of thermal resistance, compactness, and efficiency [33, 34].
LHTES design using TO was introduced by the seminal work of [11], who showed that optimized fin arrangements within PCM domains could yield branching structures that significantly improved melting and solidification rates. Subsequent studies extended this framework to multi-tube configurations [12], bio-inspired fin geometries [23], and PCM-based heat sinks whose topology-optimized designs were experimentally validated [35]. More recent contributions have examined battery thermal management systems incorporating PCMs [36], and transient PCM-integrated heat sinks [37]. Furthermore, [38] applied TO to HCM/PCM composites with variable HCM volume fraction under natural convection to optimize charging behavior. Despite these advances, most TO studies select PCM and HCM material properties a priori. [39] showed that this sequential choice of materials and geometry is suboptimal, and demonstrated that material selection can strongly influence the optimal fin topology . However, their approach was restricted to a small, discrete set of materials, which limits generality and prevents exploration of the broader material design space. Addressing this limitation requires frameworks that can represent material parameters in a continuous manner, enabling geometry and material to be optimized together.
@@ 2.3 Material–Geometry Co-Design
The challenge of simultaneous material–geometry optimization has a long history in mechanical design. Ashby’s charts and indices provided intuitive performance metrics for simple cases [40, 24] and remain influential in early-stage design. Yet, these methods are inadequate for multiphysics problems such as LHTES, where geometry and material choices are strongly coupled and the governing physics is nonlinear and transient [39].
To address these challenges, researchers explored smooth material functions [41], hybrid approaches such as genetic algorithms and sequential quadratic programming (SQP) [42] and mixed-discrete formulations [43]. While these methods provided valuable insights, they often suffer from high computational cost and limited scalability [7, 44]. In LHTES, similar limitations have restricted co-design to the evaluation of a handful of candidate materials [39].
Despite these advances, the reliance on discrete material sets and high computational cost continues to limit progress in material–geometry co-design. ML has recently emerged as a way to address these limitations. [10, 45] demonstrated that neural networks can serve as surrogate models [46] or design representations in topology optimization, enabling efficient exploration of high-dimensional design spaces. Building on this direction, [7] introduced the use of VAEs to embed discrete material databases into continuous and differentiable latent spaces. A VAE is a type of generative neural network that learns a compressed latent representation of input data while ensuring continuity and differentiability, allowing discrete material properties to be represented in forms amenable with gradient-based optimization [14]. The present work extends the VAE-based co-design framework to LHTES, where phase change exhibits nonlinear and transient behavior. By integrating a VAE-trained decoder with a finite element solver for phase-change heat transfer, our framework enables the simultaneous optimization of material and geometry within an end-to-end differentiable environment.
@@ 3 Proposed Method
This study addresses the co-optimization of topology and material selection for latent heat thermal energy storage (LHTES) devices. We begin the formulation assuming a prescribed design domain and boundary conditions, such as the axisymmetric annular configuration shown in Figure˜2; a common setup for thermal storage systems. Furthermore, we assume the availability of a material database containing candidate high conductivity materials (HCMs) and phase change materials (PCMs). This database includes relevant properties such as thermal conductivity, specific heat capacity, and mass density for all materials, cost per unit mass for HCMs, and latent heat of fusion and melting temperature for PCMs.
@@
Conventionally, the two material classes have distinct and complementary thermal properties. HCMs, typically metals, provide high thermal conductivity, whereas PCMs, often paraffins, offer superior thermal storage capacity through both sensible and latent heat but have poor conductivity. The HCMs function as thermal conduits to distribute heat into the PCM. This establishes a natural design trade-off between the system’s total heat storage capacity and its ability to transfer heat in and out of the system effectively.
Then, the design objective is to determine the optimal spatial distribution and selection of the HCM and PCM. Particularly, this work extends beyond prior geometric optimization studies by concurrently optimizing for the material choice for both constituents from the database. We employ gradient-based optimization to maximize the amount of energy discharged in a given period of time, subject to a cost constraint on the amount of HCM used.
The governing PDE and their finite element formulation are discussed in Sections˜3.1 and 3.4 respectively. The material distribution in the design is represented using conventional density-based methods employed in topology optimization (TO), as detailed in Section˜3.2. To make the discrete material database amenable to gradient-based optimization, we first convert it into a continuous, differentiable form using a Variational Autoencoder (VAE), as detailed in Section˜3.3. The decoder component of this model is then employed within the gradient-based optimization algorithm, as discussed in Section˜3.5. A complete summary of the optimization algorithm is provided in Section˜3.7.
@@ 3.1 Governing Equations
This study focuses on the TO of the thermal energy discharge process in a two-dimensional domain composed of an HCM and a PCM. The computational domain, $\Omega_{D}$, represents a quarter cross-section of the thermal energy storage (TES) unit (Figure˜2(b)). This domain eventually is partitioned, through TO, into two non-overlapping subdomains, $\Omega_{HCM}$ and $\Omega_{PCM}$, such that their union forms the complete design domain: $\Omega_{D}=\Omega_{HCM}\cup\Omega_{PCM}$.
The domain is initialized with a uniform temperature $T_{I}$. For the discharge process, a constant temperature $T_{d}$ is prescribed at the internal boundary $\Gamma_{d}$, simulating contact with a cold heat transfer fluid (HTF). The solidification process within the PCM is assumed to be dominated by heat conduction; neglecting convection. To model the heat transfer behavior, the transient heat conduction equation is employed. The governing equation, along with the associated boundary and initial conditions, can be expressed as:
@@
Here, $T$ is the temperature field, $\rho$ is the density, $\breve{c}$ is the effective specific heat capacity, $k_{ij}$ is the isotropic thermal conductivity tensor. The Neumann boundary condition on $\Gamma_{N}$ is considered to be zero to enforce the adiabatic and symmetry conditions, where $\mathbf{n}$ is the outward-pointing normal vector. To account for phase change, we adopt the apparent heat capacity method [11], where the material’s heat capacity $c_{p}$,is augmented with the latent heat of fusion, $L$, to arrive at the temperature-dependent effective specific heat capacity $\breve{c}(T)$:
@@
where the smooth indicator function, $\Pi(T)$, approximates a rectangular pulse (Figure˜3(b)) active within the mushy zone temperature range $T_{m},T_{m}+\Delta T]$ and is constructed as the difference between two smooth step functions [11]:

### v2

@@ 1 Introduction
The imperative for decarbonization [carnie2024decarbonising], grid stabilization [aftab2021phase], and waste-heat recovery [huang2023performance] has increased the demand for advanced energy storage technologies. Among competing technologies, latent-heat thermal energy storage (LHTES)—which stores energy in the latent heat of phase-change materials (PCMs)—has emerged as a promising candidate due to its high storage density [sharma2005latent]. The realization of next-generation LHTES systems is seen as critical, supported by advances in both materials and system design.
@@
On the materials front, advances in the discovery and design of materials have expanded the palette of energy-storage and transport materials, with improved storage capacity, thermal conductivity, and stability [khare2012selection, sharma2005latent]. In parallel, new manufacturing methods—particularly additive and hybrid manufacturing—have enabled the fabrication of complex heat-conduit geometries that can substantially enhance thermal transport within LHTES systems [freeman2023advanced]. Together, these developments create both opportunity and complexity for the practicing engineer in designing optimal LHTES systems.
@@
In particular, effectively leveraging this expanded material and geometric design space to extract optimal performing configurations presents a central challenge [chandrasekhar2022integrating]. Designers must select from a vast and diverse materials space while accounting for the system’s nonlinear, transient, and strongly coupled thermophysical behavior. Furthermore, the geometric freedom afforded by modern manufacturing surpasses the capabilities of traditional, intuition-driven design workflows. Consequently, these manual design approaches are limited in scope and frequently overlook high-performance configurations.
@@
To overcome these limitations, computational methods offer a systematic path forward. Specifically, topology optimization (TO) [sigmund2013_TO_review, duysinx1998topology] and machine learning (ML) [woldseth2022TO_ML_review] have emerged as powerful tools for automating design exploration. TO employs gradient-based optimization to navigate complex design landscapes, systematically discovering novel geometries and material layouts to improve performance while respecting governing physics and design constraints. Complementing this physics-driven approach, ML methods excel at discovering, organizing, and navigating expansive material databases to inform design selection [chandrasekhar2022integrating].
@@
The integration of TO and ML is well-positioned to improve the performance of LHTES devices. A critical step toward this goal is to address the strong coupling between geometry and material properties, which necessitates their simultaneous optimization. Hitherto, applications of TO have largely ignored this coupling, focusing primarily on geometric optimization while fixing the material choice a priori [pizzolato2017LHTES, pizzolato2017TO_PCM, wang2023topology]. For instance, consider a vertical tube carrying a cold heat-transfer fluid (HTF), which is surrounded by both a low-conductivity PCM and a high-conductivity material (HCM) as shown in fig.˜1(a). In traditional approaches, the PCM and HCM are first selected a priori from a material database (fig.˜1 (b)), and their spatial distribution (fig.˜1(c)) is subsequently optimized using the prescribed material properties. Consequently, the more challenging problem of co-designing geometry and material has received far less attention.
@@
A primary barrier to this simultaneous optimization is that material selection is inherently discrete, whereas gradient-based topology optimization requires a differentiable design space. This fundamental incompatibility between discrete material choice and continuous geometric optimization makes true co-design—the joint optimization of material and geometry—exceptionally challenging.
@@
This paper directly addresses this methodological gap. Building upon recent work [chandrasekhar2022integrating], we introduce an ML-enabled framework that bridges the discrete and continuous domains. Our approach employs a variational autoencoder (VAE) [kingma2013auto] to embed discrete material candidates into a continuous and differentiable latent space. This learned material representation is then integrated directly into a density-based topology optimization formulation, enabling the simultaneous co-design of material and geometry within a single, unified framework. We demonstrate that this co-design methodology yields LHTES devices with performance superior to both conventional baseline designs and sequential optimization workflows.
@@ 2 Related Work
This section surveys the state of the art in the design of LHTES devices. The survey is structured around the central contribution of this paper: the concurrent material selection and topology optimization through an ML augmented framework for the design of LHTES devices. We organize the review into three subsections. The first subsection reviews TES devices and the materials commonly employed therein. The second subsection surveys the use of TO in thermal management, in particular LHTES devices. The third subsection examines the co-design of material and geometry, outlining the limitations of conventional approaches and recent developments that incorporate machine learning.
@@ 2.1 Thermal Energy Storage
The rapid growth of renewable energy and the increasing variability of supply and demand highlight the need for next generation energy storage systems [sarbu2018comprehensive]. A wide range of storage technologies has been developed to address these challenges, including electrochemical systems [yang2011electrochemical], mechanical storage [alami2020mechanical], and TES [sarbu2018comprehensive], each with distinct operating principles, efficiency levels, costs, and environmental impacts [gur2018review, rahman2022recent]. Among these, TES plays a vital role in applications involving heat management and the integration of renewable energy [rehman2015pumped]. TES systems are typically classified into three categories: sensible heat, latent heat, and thermochemical storage [sarbu2018comprehensive]. Sensible storage, while widely deployed in hot water tanks and packed-bed systems, generally exhibits lower storage densities compared to latent and thermochemical TES types [sarbu2018comprehensive, zhang2016thermal]. Furthermore, thermochemical storage can achieve high energy density, but it relies on reversible chemical reactions that require high material stability and introduce significant complexity in terms of kinetics and long-term durability [bauer2021fundamentals, gur2018review]. Latent heat storage combines high storage density with nearly isothermal charging and discharging, making it well suited for renewable integration and thermal management applications [sarbu2018comprehensive, pizzolato2017TO_PCM, tian2021bionic].
LHTES systems rely on PCMs as the medium for storing and releasing energy [pizzolato2017LHTES]. The performance of an LHTES device, therefore, depends critically on the type of PCM employed and how it is integrated with HCMs for heat transfer. Broadly, PCMs are categorized into three groups: organics, inorganics, and eutectics [sarbu2018comprehensive]. To begin with, organic PCMs , such as paraffins and fatty acids, are widely used because they are chemically stable, non-corrosive, and readily available [sarbu2018comprehensive]. However, their low thermal conductivity often limits heat transfer rates, which necessitates the inclusion of HCMs such as copper or aluminum to improve performance [Ashby1993MaterialSelection, mehmood2018materialSelectionMEMSAshby, pizzolato2017TO_PCM]. In contrast, inorganic PCMs, particularly salt hydrates, offer higher volumetric storage capacity and thermal conductivity. Yet, they are frequently affected by drawbacks such as supercooling, phase segregation, and material compatibility issues that hinder long-term reliability [sarbu2018comprehensive]. Alternatively, eutectic PCMs combine two or more components in fixed ratios, producing well-defined melting points tailored to specific operating conditions [sarbu2018comprehensive]. This makes them suitable for applications that require precise temperature control. This diversity of PCM options and associated trade-offs motivates the need for advanced design approaches; the next subsection reviews the use of TO as a methodology for designing efficient LHTES systems.
@@ 2.2 Topology Optimization of Thermal Energy Storage Devices
TO has transformed the design of structural [sigmund2013_TO_review], fluid [borrvall2003topology, padhy2024fluto, padhy2025toflux], photonics [christiansen2021inverse, padhy2024photos], and thermal systems [tian2004effects], enabling automated generation of high-performance geometries that outperform traditional, manually designed configurations. Its relevance has increased further with the advent of additive manufacturing, which enables the fabrication of the complex, non-intuitive geometries produced through optimization studies [ge2020additive]. In thermal management applications, TO has been employed to enhance the performance of heat sinks, heat exchangers, and cooling plates, producing designs that consistently outperform conventional fin-and-tube solutions in terms of thermal resistance, compactness, and efficiency [mo2021topology, careri2023additive].
LHTES design using TO was introduced by the seminal work of [pizzolato2017LHTES], who showed that optimized fin arrangements within PCM domains could yield branching structures that significantly improved melting and solidification rates. Subsequent studies extended this framework to multi-tube configurations [pizzolato2017TO_PCM], bio-inspired fin geometries [tian2021bionic], and PCM-based heat sinks whose topology-optimized designs were experimentally validated [see2022experimental]. More recent contributions have examined battery thermal management systems incorporating PCMs [mo2021topologyPCMCoolingPlates], and transient PCM-integrated heat sinks [iradukunda2020transient]. Furthermore, [xiaoping2023LHTES] applied TO to HCM/PCM composites with variable HCM volume fraction under natural convection to optimize charging behavior. Despite these advances, most TO studies select material properties for PCM and HCM a priori. [pizzolato2020LHTES] showed that this sequential choice of materials and geometry is suboptimal, and demonstrated that material selection can strongly influence the optimal fin topology. However, their approach was restricted to a small, discrete set of materials, thereby limiting generality and preventing exploration of the broader material design space. Addressing this limitation requires frameworks that can represent material parameters in a continuous manner, enabling geometry and material to be optimized together.
@@ 2.3 Material–Geometry Co-Design
The challenge of simultaneous material–geometry optimization has a long history in mechanical design. Ashby’s charts and indices provided intuitive performance metrics for simple cases [ashby2011MaterialSelection, Ashby1993MaterialSelection] and remain influential in early-stage design. Yet, these methods are inadequate for multiphysics problems such as LHTES, where geometry and material choices are strongly coupled, and the governing physics is nonlinear and transient [pizzolato2020LHTES].
To address these challenges, researchers explored smooth material functions [ananthasuresh2003coGeomMaterial], hybrid approaches such as genetic algorithms and sequential quadratic programming (SQP) [roy2021hybrid] and mixed-discrete formulations [achleitner2022materialSelectionCompliant]. While these methods provide valuable insights, they often incur high computational costs and limited scalability [chandrasekhar2022integrating, nocedal1999Optimization]. In LHTES, similar limitations have restricted co-design to the evaluation of a handful of candidate materials [pizzolato2020LHTES].
Despite these advances, reliance on discrete material sets and the high computational cost continue to limit progress in material–geometry co-design. ML has recently emerged as a way to address these limitations. [woldseth2022TO_ML_review, chandrasekhar2021tounn] demonstrated that neural networks can serve as surrogate models [padhy2025voroto] or design representations in topology optimization, enabling efficient exploration of high-dimensional design spaces. Building on this direction, [chandrasekhar2022integrating] introduced the use of VAEs to embed discrete material databases into continuous and differentiable latent spaces. A VAE is a type of generative neural network that learns a compressed latent representation of input data while ensuring continuity and differentiability, allowing discrete material properties to be represented in forms amenable with gradient-based optimization [kingma2013auto]. The present work extends the VAE-based co-design framework to LHTES, where phase change exhibits nonlinear and transient behavior. By integrating a VAE-trained decoder with a finite element solver for phase-change heat transfer, our framework enables the simultaneous optimization of material and geometry within an end-to-end differentiable environment.
@@ 3 Proposed Method
This study addresses the co-optimization of topology and material selection for latent heat thermal energy storage (LHTES) devices. We begin the formulation by assuming a prescribed design domain and boundary conditions, such as the axisymmetric annular configuration shown in Figure˜2, which is typical of thermal storage systems. Furthermore, we assume the availability of a material database containing candidate high-conductivity materials (HCMs) and phase change materials (PCMs). This database includes relevant properties such as thermal conductivity, specific heat capacity, and mass density for all materials, cost per unit mass for HCMs, and latent heat of fusion and melting temperature for PCMs.
@@
Conventionally, the two material classes have distinct and complementary thermal properties. HCMs, typically metals, exhibit high thermal conductivity, whereas PCMs, often paraffins, offer superior thermal storage capacity through both sensible and latent heat but have low thermal conductivity. The HCMs function as thermal conduits to distribute heat into the PCM. This establishes a natural trade-off between the system’s total heat-storage capacity and its ability to transfer heat effectively into and out of the system.
The design objective is then to determine the optimal spatial distribution and selection of HCM and PCM. In particular, this work extends prior geometric optimization studies by concurrently optimizing the material selection for both constituents from the database. We employ gradient-based optimization to maximize the energy discharged over a given period, subject to a cost constraint on HCM use.
The governing PDE and its finite element formulation are discussed in Sections˜3.1 and 3.4 respectively. The material distribution in the design is represented using conventional density-based methods employed in topology optimization (TO), as detailed in Section˜3.2. To make the discrete material database amenable to gradient-based optimization, we first convert it into a continuous, differentiable form using a Variational Autoencoder (VAE), as detailed in Section˜3.3. The decoder component of this model is then employed within the gradient-based optimization algorithm, as discussed in Section˜3.5. A complete summary of the optimization algorithm is provided in Section˜3.7.
@@ 3.1 Governing Equations
This study focuses on the topology optimization of the thermal energy discharge process in a two-dimensional domain composed of an HCM and a PCM. The computational domain, $\Omega_{D}$, represents a quarter cross-section of the thermal energy storage (TES) unit (Figure˜2(b)). This domain eventually is partitioned, through TO, into two non-overlapping subdomains, $\Omega_{HCM}$ and $\Omega_{PCM}$, such that their union forms the complete design domain: $\Omega_{D}=\Omega_{HCM}\cup\Omega_{PCM}$.
The domain is initialized with a uniform temperature $T_{I}$. For the discharge process, a constant temperature $T_{d}$ is prescribed on the internal boundary $\Gamma_{d}$, simulating contact with a cold heat-transfer fluid (HTF). The solidification process within the PCM is assumed to be dominated by heat conduction, with convection neglected. While natural convection can enhance heat transfer rates in the liquid phase, particularly during long durations [wang2023topology], we adopt this conduction-dominated model to establish the foundational framework for simultaneous material-geometry co-design of LHTES systems. Thus, to model the heat transfer behavior, the transient heat conduction equation is employed. The governing equation, along with the associated boundary and initial conditions, can be expressed as:
@@
Here, $T$ is the temperature field, $\rho$ is the density, $\breve{c}$ is the effective specific heat capacity, $k_{ij}$ is the isotropic thermal conductivity tensor. The Neumann boundary condition on $\Gamma_{N}$ is considered to be zero to enforce the adiabatic and symmetry conditions, where $\mathbf{n}$ is the outward-pointing normal vector. To account for phase change, we adopt the apparent heat capacity method [pizzolato2017LHTES], where the material’s heat capacity $c_{p}$, is augmented with the latent heat of fusion, $L$, to arrive at the temperature-dependent effective specific heat capacity $\breve{c}(T)$:
@@
where the smooth indicator function, $\Pi(T)$, approximates a rectangular pulse (Figure˜3(b)) active within the mushy zone temperature range $T_{m},T_{m}+\Delta T]$ and is constructed as the difference between two smooth step functions [pizzolato2017LHTES]: